When the reddish-brown mercury(II) oxide (HgO) is heated, it decomposes to its elements, liquid mercury metal and oxygen gas. If 2.63 g HgO is decomposed to Hg, calculate the mass of the pure Hg metal produced.
Answer #1

The equation for the decomposition of HgO is:

2 HgO → 2 Hg + O2

From the equation, we see that 2 moles of HgO produce 2 moles of Hg metal, i.e moles of HG formed = moles of HgO reacted.

Moles of HgO = mass of HgO / molar mass of HgO = 2.63 / 216.6 = 0.012 moles

Mass of Hg formed = molar mass of Hg * Moles of Hg formed = 200.6 * 0.012 = 2.41 g